OpenStudy (anonymous):

Which measurement is a possible side length for a triangle with two sides measuring 9 cm and 4 cm?

OpenStudy (anonymous):

So let's draw a triangle with sides 9 and 4. 9^2 + 4^2 = c^2 81+16 =c^2 so the square root of 97 is a possible length.

OpenStudy (anonymous):

thank you <3

OpenStudy (anonymous):

So \[\sqrt{97}\] = c, the final, unkown length of this triangle. We did this using the Pythagorean theorem. \[a ^{2} + b^{2} = c^{2}\]

OpenStudy (anonymous):

Remember to include units. The cm in your final answer.

OpenStudy (anonymous):

Thank you, it all makes sense to me now ^^

OpenStudy (radar):

Note it did not state that it was a right triangle, it could be an obtuse triangle or an acute triangle, so the 3rd side could have a range of almost 13 down top almost 5 when the included angle ranges to almost 180 to 0 degrees